# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
O.1 Experimental Setup: Gaussian Sliced Wasserstein Mechanisms for Static Privatization
@@
O.1.1 Adult Census (race as secret)
@@
$S\in\mathcal{S}$
$=\{\text{White},\text{Black},\text{Asian-Pac-Islander},\text{Amer-Indian-Eskimo},\text{Other}\},$
@@
O.1.2 Cleveland Heart Disease (sex as secret)
@@
$\mathcal{Q}=\{(s_{i},s_{j}):s_{i},s_{j}\in\mathcal{S},\,s_{i}\neq s_{j}\}=\{(\text{female},\text{male}),(\text{male},\text{female})\},$
@@
O.1.3 Student Performance (grade group as secret)
@@
O.2 SRPP- ms-SRPP-SGD with Gradient Clipping
For our CIFAR-10 experiments (Figures 3(i)-3(l)) in Section 6, we use the ResNet22, adapted for compatibility with Opacus [yousefpour2021opacus] by following the architecture of [xiao2023geometry],
@@
O.3 Experiment Setup
Now, we provide detailed specifications for our SRPP-SGD and ms-SRPP-SGD experiments (Figures 3(i)-3(l)) in Section 6 of the main text.
@@
Privacy accounting: SRPP-SGD, ms-SRPP-SGD, and group-DP-SGD.
@@
(2) ms-SRPP-SGD (expected $K^{2}$ via hypergeometric model): We model the number of differing examples $K_{t}$ in each batch as a random variable following a hypergeometric distribution: $K_{t}\sim\mathrm{Hypergeometric}(N,\Delta,L)$ under fixed-size sampling without replacement. For this distribution,
@@
(3) group-DP-SGD (baseline): As a baseline, we also implement group-DP-SGD using the standard DP-SGD accountant [Abadi2016] with group size $G=\lceil\mathbb{E}_{\eta}[K_{t}^{2}]^{1/2}\rceil$ (derived from the hypergeometric model). The DP-SGD accountant computes a per-step privacy loss, which we then convert to group DP by scaling with the group size. Appendix C provides detailed discussions of the relationship between SRPP/ms-SRPP-SGD and group DP-SGD.
In the experiments for Figures 3(i) and 3(j), we sweep the privacy budget $\epsilon$ over the range $\{$ 80, 70, 65, 60, 55, 50, 45, 40, 35, 30, 25, 20, 18, 14, 12, 10, 8, 4, 2, 1, 0.5 }\} and report test accuracy at each $\epsilon$. We evaluate two sampling rates corresponding to batch sizes $L=512$ (sampling rate $q\approx 0.01$) and $L=1024$ (sampling rate $q\approx 0.02$) on the full 50,000-example training set. Training uses 40 epochs for $L=512$ and 40 epochs for $L=1024$, with clipping norm $C=4.0$ (for $L=512$) or $C=5.0$ (for $L=1024$), learning rate $\eta_{0}=0.2$ (for $L=512$) or $\eta_{0}\approx 0.14$ (for $L=1024$) with cosine decay and 5% warmup, momentum $0.9$, and weight decay $5\times 10^{-4}$. All privacy accounting uses Rényi order $\alpha=16$ and add-remove adjacency.

### v3

@@ Appendix H Minimality and Attainability of HUC
W.1 Experimental Setup: Gaussian Sliced Wasserstein Mechanisms for Static Privatization
@@
W.1.1 Adult Census (race as secret)
@@
$S\in\mathcal{S}$ $=\{\text{White},\text{Black},\text{Asian-Pac-Islander},$
$\text{Amer-Indian-Eskimo},\text{Other}\},$
@@
W.1.2 Cleveland Heart Disease (sex as secret)
@@
$\mathcal{Q}$ $=\{(s_{i},s_{j}):s_{i},s_{j}\in\mathcal{S},\,s_{i}\neq s_{j}\}$
$=\{(\text{female},\text{male}),(\text{male},\text{female})\},$
@@
W.1.3 Student Performance (grade group as secret)
@@
W.2 SRPP- sa-SRPP-SGD with Gradient Clipping
For our CIFAR-10 experiments (Figures 2(i)-2(l)) in Section 6, we use the ResNet22, adapted for compatibility with Opacus [yousefpour2021opacus] by following the architecture of [xiao2023geometry],
@@
W.3 Experiment Setup
Now, we provide detailed specifications for our SRPP-SGD and sa-SRPP-SGD experiments (Figures 2(i)-2(l)) in Section 6 of the main text.
@@
Privacy accounting: SRPP-SGD, sa-SRPP-SGD, and group-DP-SGD.
@@
(2) sa-SRPP-SGD (expected $K^{2}$ via hypergeometric model): We model the number of differing examples $K_{t}$ in each batch as a random variable following a hypergeometric distribution: $K_{t}\sim\mathrm{Hypergeometric}(N,\Delta,L)$ under fixed-size sampling without replacement. For this distribution,
@@
(3) group-DP-SGD (baseline): As a baseline, we also implement group-DP-SGD using the standard DP-SGD accountant [Abadi2016] with group size $G=\lceil\mathbb{E}_{\eta}[K_{t}^{2}]^{1/2}\rceil$ (derived from the hypergeometric model). The DP-SGD accountant computes a per-step privacy loss, which we then convert to group DP by scaling with the group size. Appendix G provides detailed discussions of the relationship between SRPP/sa-SRPP-SGD and group DP-SGD.
In the experiments for Figures 2(i) and 2(j), we sweep the privacy budget $\epsilon$ over the range $\{$ 80, 70, 65, 60, 55, 50, 45, 40, 35, 30, 25, 20, 18, 14, 12, 10, 8, 4, 2, 1, 0.5 }\} and report test accuracy at each $\epsilon$. We evaluate two sampling rates corresponding to batch sizes $L=512$ (sampling rate $q\approx 0.01$) and $L=1024$ (sampling rate $q\approx 0.02$) on the full 50,000-example training set. Training uses 40 epochs for $L=512$ and 40 epochs for $L=1024$, with clipping norm $C=4.0$ (for $L=512$) or $C=5.0$ (for $L=1024$), learning rate $\eta_{0}=0.2$ (for $L=512$) or $\eta_{0}\approx 0.14$ (for $L=1024$) with cosine decay and 5% warmup, momentum $0.9$, and weight decay $5\times 10^{-4}$. All privacy accounting uses Rényi order $\alpha=16$ and add-remove adjacency.